dx=(3y2+4yz)dy+(2yz+y2)dz
1
Expert's answer
2020-10-20T13:59:30-0400

"dx=(3y^2+4yz)dy+ (2yz+y^2)dz"


"\int dx=\int(3y^2+4yz)dy+\int (2yz+y^2)dz"


"x=\frac{3y^{3}}{3}+\frac{4y^{2}z}{2}+\frac{2yz^{2}}{2}+y^{2}z+c"


"x=y^{3}+2y^{2}z+yz^{2}+y^{2}z+c"